OpenStudy (anonymous):

Please help!! What is f(x)/ g(x) when f(x) = 6x^3 - 19x^2 + 16x - 4 and g(x) = x - 2

OpenStudy (anonymous):

\(\huge \frac{f(x)}{g(x)}=\frac{6x^3 - 19x^2 + 16x - 4}{x-2}\\\huge =\frac{(x-2)(3x-2)(2x-1)}{x-2} \\\huge=(3x-2)(2x-1)\)

OpenStudy (anonymous):

my choices are 1) 6x^2 - 7x + 2 2) 3x^2 - 9x + 8 3) 6x^2 - 7x + 2 -8/(x-2) 4) 3x^2 - 9x + 8 -8/(x-2)